For a one-tailed test (lower tail) at 93.7% confidence, Z =

a)

-1.86

b)

-1.53

c)

-1.96

d)

-1.645
Answer #1

solution:

93.7% confidence level the z is ,

  = 1 - 93.7% = 1 - 0.937= 0.063

/ 2 = 0.063 / 2 = 0.0315

Z/2 = Z0.0315 = -1.86